OpenStudy (anonymous):

What is the diameter of a circle that has an area of 452.16 mm 2? 144 mm 12 mm 24 mm 6 mm

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

yep

OpenStudy (skullpatrol):

$$\Huge D=2*r$$

OpenStudy (skullpatrol):

$$\Huge A=\pi r^2$$

OpenStudy (anonymous):

@skullpatrol is it 24??

OpenStudy (anonymous):

so 144 is not the answer

OpenStudy (anonymous):

since Diameter = square root of (452.16/pi)*2 = 23.99391574 or about 24 mm

OpenStudy (anonymous):

@parisgirl1515 no lol

OpenStudy (anonymous):

ok